OpenStudy (anonymous):

Given the equation, 2|x + 4| = 8, the isolated absolute value equation would be |x + 4| = 4.

OpenStudy (kirbykirby):

\[ |x+4| = \begin{cases} -(x+4) & \text{if } (x+4) < 0 \\ x+4 & \text{if } (x+4) \ge 0 \end{cases} \]

OpenStudy (kirbykirby):

so basically u solve both -(x+4)=4 (x+4)=4